2, It is known that the quantitative relationship between the dependent variable Y and the independent variables X and 2 is: A. Make a Table and draw the graph of the relationship between Y and X when the numerical value of Z is: B. Make a Table and draw the graph of the relationship between Y and Z when the numerical value of X is: C. Make a Table of the value of the rate of change of Y with respect to change of X for values of x between negative five and positive five. 3. It is known that the relationship between the dependent variable Y and the independent variables X and Z is A. Make a Table and draw a graph of the relationship between Y and z when the value of X is: b.X1
Answer #1

Solution:

Z =0 Y = X^2
Z =-1 Y = X^2 - 2
Y = X^2
X -3 -2 -1 0 1 2 3
Y 9 4 1 0 1 4 9

39 ろ 24 2 3 2 2 2 2

7 3 2 2 2 2 2 2 2 3. 3

lo) 9 χ--5 x Z. 4 5 -20246 3 -2 3 2 2

Z = seq(-5,5,by = 1)
> Y = 25 + 2*Z
> plot(Z,Y, type = 'l')
> abline(v=0)

4 2 2 -4

t= x2 + 2.2 ICJ charge of t umtx 3 10